(no subject)

Emmanuel Monnier (monnier@hep.uchicago.edu)
Fri, 18 Jun 1999 15:51:22 -0500 (CDT)

Shift Summary for OWL shift, June 18, 1999

On Shift: EM, JH

Hours of beam delivered at all intensities: 6
Hours of beam delivered at nominal intensity: 6
Hours of E832 data (special runs not included): 6
Hours of special runs (laser scans, muon runs...): 1.5

Accounting of lost time: Radiation survey started at 1:00 pm then
vacuum tripped.

========================================================================
Runs ended this shift:
Run 13644 - 15M events - BEAM_832 (23K vac00)
Run 13645 - Laser scan (with low intensity muons for rad survey)
Run 13647 - Laser scan (no muons)

Tapes brought to FCC this shift: QKC QKS QKK016-018, QKB006
(including tapes for online split)

========================================================================
Accesses during shift: Radiation survey crew, Doug, Jason H, Elizabeth,
Miguel

========================================================================
Overview of shift:
Smooth E832 data taking till the survey crew started their work
then an hour later the vacuum tripped and at the end of the shift
we were still trying to recover the vacuum.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
13644 10:00 To get the BOA on epicure two scalers were modified
Acc mu is now Acc90 in-time (nm4sclr 1)
mu2*mu3 is now Acc90 out of time (nm4sclr 2)
The BOA = ratio nm4sclr 1 / nm4sclr 2
13644 10:41 DYC 53 full kfrend restart
11:10 T5 cannot be changed so triggers between T4 and T5
will be vetoed. For that, a cable has been installed
between RR185 and RR164 (cable installed during
DA stopped at 13:30)
13644 13:00 Run stopped. Hadron run stopped for radiation
survey.
13645 13:30 Laser scan started

13645 14:00 Vacuum tripped.
13647 14:45 Laser scan started

15:00 Eizabeth, miguel, jason went on access.
15:30 still not recovered.
15:30 Jason swapped contoller for TE Etotal crate
Miguel changed DC1 threshold back to 40 mV and
checked DC gas flow pressure
13648 15:40 Cosmics run started
15:35 Analysis magnet ramped down

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift